What is the dimensions of a rectangle with a perimeter of 56 cm.if its width is 75% of its lenght?
Answer by Alan3354(69443)

What is the dimensions of a rectangle with a perimeter of 56 cm.if its width is 75% of its lenght [sic]?
======================
2L + 2W = 56
L + W = 28
W = 0.75L
------------
L + 0.75L = 28
etc